# A New Approach to Backtracking Counterfactual Explanations: A Causal Framework for Efficient Model Interpretability

### v1

@@ 8.1 Simulation Setup
To evaluate the proposed method, we adopt the experiment in causal algorithmic recourse paper (Karimi et al., 2021), as it serves as a critical baseline for comparison. Since the primary focus is on assessing the interpretability of the proposed method, we use a simple model for $h(\cdot)$. This ensures that the exact solutions to the optimization problem can be computed and aligned with our intuitive understanding of the task.
@@
Our Method ($\lambda=1$) female 24 \$4087 33.0
Our Method ($\lambda=1.2$) female 24 \$3736 33.3
@@
Method Gender Age Loan amount Duration
Original (High-risk) male 27 \$14027 60
Our Method ($\lambda=1$) male 27 \$13686 36.9
Our Method ($\lambda=1.2$) male 27 \$13149 37.4
Causal Algorithmic Recourse (Karimi et al., 2021) male 27 \$14027 36.6
@@ 8.3 Sensitivity Analysis
Estimating causal functions in the input graph often involves approximations, introducing potential noise. To evaluate the robustness of our method, we add zero-mean Gaussian noise with a standard deviation of 5 to the coefficients of $f_{3}(\cdot)$ and solve the optimization problem (26) with $\lambda=1.2$. The resulting counterfactuals are:
@@ 9 Conclusion
In this work, we presented a new framework for counterfactual explanations based on backtracking counterfactuals. Our approach overcomes the limitations of interventional counterfactuals by introducing an optimization problem that generates actionable and causally consistent explanations. We showed that our method unifies several existing approaches—such as counterfactual explanations, deep backtracking explanations, backtracking counterfactuals, and causal algorithmic recourse—under a single framework. Additionally, we demonstrated that our method is both easier to understand and more computationally efficient compared to causal algorithmic recourse. Through simulation experiments, we verified that the proposed method produces explanations that are more intuitive for users and more practical for real-world applications.
@@ 10 Future Work
Relaxing Assumptions for Connection with Causal Algorithmic Recourse: Our approach depends on the convexity and ANM assumptions to establish a connection with causal algorithmic recourse. Future work could investigate alternative conditions that do not require these assumptions, allowing the theory to be applied to non-linear and non-convex models frequently found in real-world applications.

### v2

@@ 8.1 Simulation Setup
To evaluate the proposed method, we adopt the experiment in the Causal Algorithmic Recourse paper (Karimi et al., 2021), as it serves as a critical baseline for comparison. Since the primary focus is on assessing the interpretability of the proposed method, we use a simple model for $h(\cdot)$. This ensures that the exact solutions to the optimization problem can be computed and aligned with our intuitive understanding of the task.
@@
Brace (Our Method, $\lambda=1$) female 24 \$4087 33.0
Brace (Our Method, $\lambda=1.2$) female 24 \$3736 33.3
@@ 8.2 Optimization Problem and Results
Method Gender Age Loan amount Duration
Original (High-risk) male 27 \$14027 60
Brace (Our Method, $\lambda=1$) male 27 \$13686 36.9
Brace (Our Method, $\lambda=1.2$) male 27 \$13149 37.4
Causal Algorithmic Recourse (Karimi et al., 2021) male 27 \$14027 36.6
@@ 8.3 Sensitivity Analysis
Estimating causal functions in the input graph often involves approximations, introducing potential noise. To evaluate the robustness of our method, we add zero-mean Gaussian noise with a standard deviation of 5 to the coefficients of $f_{3}(\cdot)$ and solve the optimization problem (25) with $\lambda=1.2$ for an individual with features $\mathbf{x}=(\text{female},24,\$4308,48)$. The resulting counterfactuals are:
@@ 9 Conclusion
In this work, we presented a new framework Brace for counterfactual explanations based on backtracking counterfactuals. Our approach overcomes the limitations of interventional counterfactuals by introducing an optimization problem that generates actionable and causally consistent explanations. By solving a single unified objective parameterized by $\lambda$, Brace recovers four established paradigms—classical Counterfactual Explanations ($\lambda=0$), Deep Backtracking Explanations ($\lambda\to\infty$), Backtracking Counterfactual Explanations via a specific backtracking conditional distribution, and Causal Algorithmic Recourse under a convexity assumption. Additionally, we demonstrated that our method is both easier to understand and more computationally efficient compared to causal algorithmic recourse. Through simulation experiments, we verified that the proposed method produces explanations that are more intuitive for users and more practical for real-world applications.
@@ 10 Future Work
Relaxing Assumptions for Connection with Causal Algorithmic Recourse: Our approach depends on convexity assumptions to establish a connection with causal algorithmic recourse. Future work could investigate alternative conditions that do not require these assumptions, allowing the theory to be applied to non-linear and non-convex models frequently found in real-world applications.